# High CO/H2 ratios supports an exocometary origin for a CO-rich debris disk

### v2

@@ 4.2 H2 + self-shielding of CO: likely insufficient in a primordial scenario
We have also taken other assumptions such as that the CO and H2 column densities measured by CRIRES+ are representative of the disk midplane, although both disks’ inclinations are not perfectly edge-on (Table 1). This is reasonable given that the CO in both disks peaks relatively close to the star in both systems; for example HD 131488 has a CO inner edge $\lesssim 5$ AU and is vertically resolved in recent high-resolution ALMA data from the ARKS large program (Mac Manamon et al. in press). Finally, our calculation assumes the photodissociation lifetime is representative of the survival times of CO molecules in the disk in the absence of other destruction/production mechanisms, such as gas-phase chemistry reforming CO from other species. However, this is only expected to be important in H2-rich environments (Higuchi et al., 2017; Iwasaki et al., 2023), which is unlikely to be the case here given our stringent H2 upper limits.
Overall, most of these assumptions imply that our CO shielding effect and in turn photodissociation timescales in the two systems, calculated from our line of sight column densities of CO and H2, are likely significantly overestimated. This could imply that contrary to the values we report, H2 is likely not able to provide sufficient shielding and allow CO to survive over the lifetime (age) of the systems. This would provide further support to the claim that the gas is not primordial but of second generation in these systems, adding to the new conclusive evidence presented here that the CO/H2 ratio differs significantly from protoplanetary disk gas in at least one of our two CO-rich systems. Nonetheless, accurately modelling the disk geometry, local radiation field, shielding from various molecules, outgassing rates, etc. would need to be conducted to determine a more realistic photodissociation timescale.
@@
For both disks, it also is in principle possible that the belts host a hybrid mix of primordial and secondary gas, with their protoplanetary gas having started to dissipate very recently, as suggested by Lisse et al. (2017). However, our evidence suggests that even in such a scenario, the disks are likely still dominated by second generation gas. Our non-detection of H2 in both systems is consistent with abundance ratios found in Solar System comets (Combi, 1996); supporting the hypothesis that the gas in these CO-rich systems is, like the dust, secondary/exocometary in origin rather than primordial.
Given the conservative nature of our calculation, even though our photodissociation timescale is older than the age of the system, it is unlikely that any unseen H2 is sufficient to provide adequate shielding to the CO gas on it’s own for HD 131488. This is not conclusive evidence against a primordial origin scenario and more complete modelling would need to be conducted to determine a more realistic photodissociation timescale. However, for HD 110058, the $\frac{\text{CO}}{\text{H}_{2}}$ ratio still conclusively rules out that the gas has a primordial-like composition.
@@ 4.3 Challenges remain for the second-generation model
While our results strongly point to a secondary origin for the gas at least for HD 110058, challenges remain for modelling secondary gas production and survivability in exocometary belts. Producing the observed column densities in CO-rich belts through exocometary release without shielding may require unreasonably large gas release rates to counter the destruction timescale of unshielded CO, which is $\sim$ 130 years in the ISRF (Heays et al., 2017). To produce the CO mass of $4-8\times 10^{-2}M_{\bigoplus}$ found in HD 21997 for example, Kóspál et al. (2013) estimates that approximately 6000 Hale-Bopp sized comets would need to be destroyed yearly. CI (as produced through CO photodissociation) has been posited as a shielding agent to allow for a slower CO release rate and longer accumulation timescales (Matrà et al., 2017a; Marino et al., 2020).
@@
Indeed, the molecular composition of gas in CO-rich exocometary belts also remains an open question. Constraints for the presence of molecules have been placed on $\beta$ Pictoris (Matrà et al., 2018b), 49 Ceti (Klusmeyer et al., 2021), HD 21997, HD 121617, HD 131488, and HD 131835 (Smirnov-Pinchukov et al., 2022). The upper limits found for molecular species in these works differ from the younger protoplanetary disks for which similar searches have been conducted (Smirnov-Pinchukov et al., 2022). The limits are consistent with the secondary generation scenario (Matrà et al., 2018b) but in the case of CN indicate that CO could be preferentially shielded compared to other molecules (Klusmeyer et al., 2021). Higher signal-to-noise ratio observations are likely needed to find molecules other than CO in gas rich exocometary belts. This unprecedented deep search for molecular hydrogen demonstrates the usefulness of IR instruments like CRIRES+ for constraining the composition of exocometary gas in edge-on exocometary belts. In particular, if the gas is exocometary as strongly supported by our H2 constraints, observations of IR lines in absorption and mm lines in emission can be used to probe the molecular, volatile content of exocometary material. This is of interest for systems such as HD 131488 and HD 110058 which are a few 10s of Myrs-old and are therefore in the latest stages of terrestrial planet formation when volatile delivery is most likely to occur (Morbidelli et al., 2012, and references therein).
@@ 5 Conclusions
  • We report detections of the 12CO v=2-0 rovibrational transitions in HD 131488 and HD 110058 and calculate best-fit, line of sight column densities of 18.05 cm-2 and 19.97 cm-2 respectively. We find broad agreement with previous UV observations with HST of these disks, with the exception of the kinetic temperatures, which is likely due to the lines being spectrally unresolved, and the radial velocity of HD 131488, although our radial velocity is consistent with mm observations with ALMA.
@@
  • Combined with our CO detections, we compare the $\frac{CO}{\text{H}_{2}}$ ratio lower limits of $3.09\times 10^{-5}$ (HD 131488) and $1.35\times 10^{-3}$ (HD 110058) with protoplanetary disks and conclude that the composition of HD 110058 is significantly different to primordial gas found in protoplanetary disks while that of HD 131488 remains potentiall consistent with younger disks at low probability. Further observations would be needed to confirm whether HD 131488 has a primordial or secondary composition.
  • H2 column densities are unlikely to provide sufficient shielding for CO to survive for the system age for both systems. This is because while a very conservative radial estimate leads to timescales of 174 Myr and 61 Myr for HD110058 and HD131488, these values are likely orders of magnitude lower when accounting for photodissociation in the vertical direction and by stellar photons.
  • Together, these lines of evidence imply that the gas observed in these CO-rich disks is likely second generation gas released by exocomets in the disk, rather than primordial gas leftover from their parent protoplanetary disks.
@@ Appendix A Molecfit Telluric Corrections
Figure 7: A portion of an example spectrum (blue) overlaid with the telluric fit (orange) produced by molecfit over the data from a single nodding position/day combination (the combined nodding A frames on the night of 08/03/2024). The locations of 12CO lines are marked with green dashed lines.
@@ Appendix B On CRIRES+ wavelength calibration offset between nodding positions
Our observations are taken with the 0.2” width slit on CRIRES. When reducing our data, the ESO data reduction pipeline reported that for both of our targets, the width of the slit was not fully illuminated. This has two consequences for our observations detailed in the ESO CR2RES pipeline user manual in the known issues section 7.1 on ”Superresolution”. Relevant to this study, this gives us increased resolution but at the cost of the default wavelength solution produced by the ESO pipeline being unreliable between nodding frames. To account for this, we take the combined spectra for each nodding position and separately feed each one through molecfit to acquire a final wavelength solution fit to the tellurics in the data. For each star this means we fit 8 spectra for 12CO (2 days, 2 nodding positions, and 2 sub-orders) and 4 spectra for H2 (2 days and 2 nodding positions). When presenting the results in this paper, we show plots of the median of these spectra. Any combination or modelling of spectra with an incorrect wavelength solution would distort the line shapes and positions, diminishing the accuracy of our results. The degree to which this is an issue is shown in Figures 8 and 9 which show the pre-corrected wavelength solution can err by as much as a pixel. The higher post-correction correlation in Figure 9 demonstrates numerically that the alignment of the spectra is improved.
@@
Figure 8: 12CO lines before and after superresolution correction. In red, the nodding A frames should align with the nodding B frames as these observations were taken on the same day if the wavelength solution is accurate.
@@
Figure 9: Pearson correlations of the uncorrected (blue) and post-correction (orange) 12CO lines from Figure 8.
Appendix C MCMC Posterior Distributions
@@
In Figures 10, and 11 we show the posterior distributions from the MCMC simulations for H2. Given that H2 is not detected in either system, the 2d corner plots show upper limits for the H22 column density and unconstrained temperature as expected for a non-detection. Higher temperatures allow for larger column densities as the thermal broadening produces shallower, wider lines which are more difficult to detect than the narrower, deeper lines produced by cold gas.
Figure 10: Posterior probability distributions of excitation temperature and column density from the 2223.29 nm rovibrational H2 absorption model fit to the HD 131488 data. Marginalised distributions for each parameter are displayed on the diagonal.
Figure 11: Posterior probability distributions of excitation temperature and column density from the 2223.29 nm rovibrational H2 absorption model fit to the HD 110058 data. Marginalised distributions for each parameter are displayed on the diagonal.
Figure 12: Posterior probability distributions of excitation temperature, kinetic temperature, radial velocity, and column density from the for the 12CO absorption model fit to the HD 131488 data. Marginalised distributions for each parameter are displayed on the diagonal.
Figure 13: Posterior probability distributions of excitation temperature, kinetic temperature, radial velocity, and column density from the for the 12CO absorption model fit to the HD 110058 data. Marginalised distributions for each parameter are displayed on the diagonal.
@@ Appendix D CO/H2 Ratios for Protoplanetary and Exocometary Belt Gas
(1) Zhang et al. (2020); (2) Robitaille et al. (2007); (3) Artur de la Villarmois et al. (2018); (4) Zhang et al. (2019); (5) Wichittanakom et al. (2020); (6) Anderson et al. (2019); (7) Kama et al. (2016); (8) Bruderer et al. (2012); (9) Pineda et al. (2019); (10) Asensio-Torres et al. (2021); (11) Schneider et al. (2018); (12) Natta et al. (2006); (13) Mannings & Sargent (2000); (14) Andrews et al. (2009); (15) Macías et al. (2018); (16) Müller et al. (2011); (17) Pecaut et al. (2012); (18) Merín et al. (2004); (19) Kraus & Hillenbrand (2009); (20) Simon et al. (2000); (21) Vioque et al. (2018); (22) Makarov (2007); (23) Neuhäuser et al. (2000); (24) Torres et al. (2008); (25) This work; (26) Pecaut et al. (2012); (27) Pecaut et al. (2012);(28) Rhee et al. (2007); (29) Mawet et al. (2012); (30) Trapman et al. (2025); (31) Bergin & Williams (2018); (32) Siess et al. (2000); (33) D’Antona & Mazzitelli (1997); (34) Luhman & Rieke (1999), (35) Zaire et al. (2024) (36) Lee et al. (2024); (37) Lecavelier des Etangs et al. (2001)

### v3

@@ 4.2 H2 + self-shielding of CO: likely insufficient in a primordial scenario
ALMA images of the same CO gas (MacManamon et al. in prep.) show the disk has a vertical height above the midplane of $\sim$ 20 au (where the FWHM of the resolution element for this image was 11 au) and a radial extent of $\sim$ 180 au. The shielding factors of Visser et al. (2009) assume an isotropic radiation field and a spherical cloud shielding a CO molecule in the centre of that cloud. Using the disk’s radial column density as the radius of a spherical cloud, we obtain an upper limit of the photodissociation timescale of 55.4 Myr. If, on the other hand, we take the ratio of the height to radius of the disk we can scale the radial column density to estimate a vertical column density. Calculating the shielding provided by a spherical cloud with radius equal to the vertical half height we obtain a lower limit of the photodissociation timescale of 0.11 Myr. The true value is therefore likely between 0.11 and 55.4 Myr; given the disk age of 16 Myr, we cannot formally rule out a primordial origin scenario for HD 131488.
@@
Another assumption affecting the calculation is that the CO and H2 column densities measured by CRIRES+ are representative of the disk midplane and thus the bulk of the gas, although both disks’ inclinations are not perfectly edge-on (Table 1). However, the CO in both disks peaks relatively close to the star in both systems, meaning that the bulk of the gas is in a narrow vertical distribution. HD 131488 has a CO inner edge $\lesssim 5$ AU and is vertically resolved in recent ALMA data (Mac Manamon et al. in prep.). Finally, our calculation assumes the photodissociation lifetime is representative of the survival times of CO molecules in the disk in the absence of other destruction/production mechanisms, such as gas-phase chemistry reforming CO from other species. However, this is only expected to be important in H2-rich environments (Higuchi et al. 2017; Iwasaki et al. 2023), which is unlikely to be the case here given our stringent H2 upper limits.
Overall, most of these assumptions imply that the CO shielding and in turn photodissociation timescales in HD 131488 calculated from our line of sight column densities of CO and H2, is likely significantly overestimated. Although our quantitative estimate is formally not conclusive, H2 is likely not able to provide sufficient shielding and allow CO to survive over the lifetime (age) of the systems. This would provide further support to the claim that the gas is not primordial but of second generation in HD 131488 as well as in HD 110058, adding to the new conclusive evidence presented here that the CO/H2 ratio differs significantly from protoplanetary disk gas in at least one of our two CO-rich systems. Nevertheless, accurately modelling HD 131488’s geometry, local radiation field, shielding from various molecules, outgassing rates, etc. would need to be conducted to determine a more realistic photodissociation timescale.
@@
For both disks, it is in principle possible that the belts host a hybrid mix of primordial and secondary gas, with their protoplanetary gas having started to dissipate very recently, as suggested by Lisse et al. (2017). However, our evidence suggests that even in such a scenario, at least HD 110058 is likely still dominated by second generation gas. Our non-detection of H2 in both systems is consistent with abundance ratios found in Solar System comets (Combi 1996); supporting the hypothesis that the gas in these CO-rich systems is, like the dust, secondary/exocometary in origin rather than primordial.
@@ 4.3 Challenges remain for the second-generation model
While our results strongly point to a secondary origin for the gas at least for HD 110058, challenges remain for modelling secondary gas production and survivability in exocometary belts. Producing the observed column densities in CO-rich belts through exocometary release without shielding may require unreasonably large gas release rates to counter the destruction timescale of unshielded CO, which is $\sim$ 130 years in the ISRF (Heays et al. 2017). To produce the CO mass of $4-8\times 10^{-2}M_{\bigoplus}$ found in HD 21997 for example, Kóspál et al. (2013) estimates that approximately 6000 Hale-Bopp sized comets would need to be destroyed yearly. CI (as produced through CO photodissociation) has been posited as a shielding agent to allow for a slower CO release rate and longer accumulation timescales (Matrà et al. 2017a; Marino et al. 2020).
@@
Indeed, the molecular composition of gas in CO-rich exocometary belts also remains an open question. Constraints for the presence of molecules have been placed on $\beta$ Pictoris (Matrà et al. 2018b), 49 Ceti (Klusmeyer et al. 2021), HD 21997, HD 121617, HD 131488, and HD 131835 (Smirnov-Pinchukov et al. 2022). The upper limits found for molecular species in these works differ from the younger protoplanetary disks for which similar searches have been conducted (Smirnov-Pinchukov et al. 2022). The limits are consistent with the secondary generation scenario (Matrà et al. 2018b) but in the case of CN indicate that CO could be preferentially shielded compared to other molecules (Klusmeyer et al. 2021). Higher signal-to-noise ratio observations are likely needed to find molecules other than CO in gas rich exocometary belts. This unprecedented deep search for molecular hydrogen demonstrates the usefulness of IR instruments like CRIRES+ for constraining the composition of exocometary gas in edge-on exocometary belts. In particular, if the gas is exocometary as strongly supported by our H2 constraints, observations of IR lines in absorption and mm lines in emission can be used to probe the molecular, volatile content of exocometary material. This is of interest for systems such as HD 131488 and HD 110058 which are a few 10s of Myrs-old and are therefore in the latest stages of terrestrial planet formation when volatile delivery to these planets is most likely to occur (Morbidelli et al. 2012, and references therein).
@@ 5 Conclusions
  • We report detections of the 12CO v=2-0 rovibrational transitions in HD 131488 and HD 110058 and calculate best-fit, line of sight log column densities of 18.05 cm-2 and 19.97 cm-2 respectively.
@@
  • Combined with our CO detections, we compare the $\frac{\text{CO}}{\text{H}_{2}}$ ratio lower limits (assuming the CO and H2 are at the same temperature) of $3.09\times 10^{-5}$ (HD 131488) and $1.35\times 10^{-3}$ (HD 110058) with protoplanetary disks and conclude that the composition of HD 110058 is significantly different to primordial gas found in protoplanetary disks while that of HD 131488 remains potentially consistent with younger disks.
  • We calculate conservative upper and lower limits for the photodissociation timescale of the CO in HD 131488, finding that we cannot formally rule out a primordial origin scenario for the gas with a simple shielding model. Given the conservative assumptions, we postulate that a more realistic shielding model is likely to rule out a primordial origin scenario. Further observations would be needed to confirm whether HD 131488 has a primordial or secondary composition.
@@ Appendix A Molecfit telluric corrections
Figure 7: A portion of an example spectrum (black) overlaid with the telluric fit (blue) produced by molecfit over the data from a single nodding position/day combination (the combined nodding A frames on the night of 08/03/2024). The locations of 12CO lines are marked with green dashed lines.
@@ Appendix B On CRIRES+ wavelength calibration offset between nodding positions
Our observations are taken with the 0.2” width slit on CRIRES. When reducing our data, the ESO data reduction pipeline reported that for both of our targets, the width of the slit was not fully illuminated. This has two consequences for our observations detailed in the ESO CR2RES pipeline user manual in the known issues section 7.1 on ”Superresolution”. Relevant to this study, this gives us increased resolution but at the cost of the default wavelength solution produced by the ESO pipeline being unreliable between nodding frames. To account for this, we take the combined spectra for each nodding position and separately feed each one through molecfit to acquire a final wavelength solution fit to the tellurics in the data. For each star this means we fit 8 spectra for 12CO (2 days, 2 nodding positions, and 2 sub-orders) and 4 spectra for H2 (2 days and 2 nodding positions). When presenting the results in this paper, we show plots of the median of these spectra. Any combination or modelling of spectra with an incorrect wavelength solution would distort the line shapes and positions, diminishing the accuracy of our results. The degree to which this is an issue is shown in Figures 8 and 9 which show the pre-corrected wavelength solution can err by as much as a pixel. The higher post-correction Pearson correlations of the A and B nodding frames in Figure 9 demonstrates numerically that the alignment of the spectra is improved.
@@
Figure 8: 12CO lines before and after superresolution correction. In red, the nodding A frames should align with the nodding B frames in blue as these observations were taken on the same day if the wavelength solution is accurate.
@@
Figure 9: Pearson correlations of the nodding A and B positions for the uncorrected (blue) and post-correction (orange) 12CO lines from Figure 8.
Appendix C MCMC posterior/prior distributions
@@
Table 4 gives the priors used in our MCMC simulations. In Figures 10, and 11 we show the posterior distributions from the MCMC simulations for H2 and these are also stated in Table 3. Given that H2 is not detected in either system, the 2d corner plots show upper limits for the H2 column density and unconstrained temperature as expected for a non-detection. Higher temperatures allow for larger column densities as the thermal broadening produces shallower, wider lines which are more difficult to detect than the narrower, deeper lines produced by cold gas.
For CO, each prior bound was selected to be much larger than the posterior CO distributions found by Brennan et al. (2024). The H2 column density priors allow the H2 to be the same as the CO or up to $10^{5}$ times larger (an order of magnitude larger than an ISM-like abundance). This allows for anything between a non-detection and line profiles much larger than those we would expect for a detection. The temperature range for H2 was chosen to be smaller to speed up convergence while still being much larger than the CO posterior temperatures. Smaller column density ranges were used in test runs and found not to significantly influence the upper limits calculated later. We ran each MCMC for 50000 steps with 100 walkers with a burn-in of the first 10% of the steps and we visually inspected the chains to ensure convergence.
Table 4: Bounds of the uniform priors for each parameter in the MCMCs.
Parameter CO H2
log(N (cm-2)) $10^{15}\rightarrow 10^{23}$ $10^{19}\rightarrow 10^{23}$
vr (HD 110058) (km s-1) $10\rightarrow 15$ CO best fit $\pm 0.5$
vr (HD 131488) km s-1 $2\rightarrow 8$ CO best fit $\pm 0.5$
[…]
@@ Appendix D CO/H2 ratios for protoplanetary and exocometary belt gas
(1) Zhang et al. (2020); (2) Robitaille et al. (2007); (3) Artur de la Villarmois et al. (2018); (4) Zhang et al. (2019); (5) Wichittanakom et al. (2020); (6) Anderson et al. (2019); (7) Kama et al. (2016); (8) Bruderer et al. (2012); (9) Pineda et al. (2019); (10) Asensio-Torres et al. (2021); (11) Schneider et al. (2018); (12) Natta et al. (2006); (13) Mannings and Sargent (2000); (14) Andrews et al. (2009); (15) Macías et al. (2018); (16) Müller et al. (2011); (17) Pecaut et al. (2012); (18) Merín et al. (2004); (19) Kraus and Hillenbrand (2009); (20) Simon et al. (2000); (21) Vioque et al. (2018); (22) Makarov (2007); (23) Neuhäuser et al. (2000); (24) Torres et al. (2008); (25) This work; (26) Pecaut et al. (2012); (27) Pecaut et al. (2012);(28) Rhee et al. (2007); (29) Mawet et al. (2012); (30) Trapman et al. (2025); (31) Bergin and Williams (2017); (32) Siess et al. (2000); (33) D’Antona and Mazzitelli (1997); (34) Luhman and Rieke (1999), (35) Zaire et al. (2024) (36) Lee et al. (2024); (37) Lecavelier des Etangs et al. (2001)